OpenStudy (anonymous):

Please help me find the function? and Explain why ? (x , y) (1 , 1) (2 , 3) (3 , 9) (4 , 27)

OpenStudy (anonymous):

This is tough to explain. I see a pattern of powers of 3. 3^0=1, 3^1=3, 3^2=9, 3^3=27 To write an expression it would be 3^x-1 = y

OpenStudy (anonymous):

This problem doesn't work like the last, because if you try it by difference there isn't a consistent one.

OpenStudy (anonymous):

\[y=3^{x-1}\]

OpenStudy (anonymous):

I'm trying to think of a systematic way to approach this problem....

OpenStudy (anonymous):

No consistent difference told me it was either a constant to the power of x

OpenStudy (anonymous):

yeah this problems confusing haha but i get what you getting at. but If this pattern were to continue, what would the 7th ?

OpenStudy (anonymous):

I just started thinking about the numbers and they were all divisible by 3.

OpenStudy (anonymous):

the 7th would be 3^(7-1). I don't have a calculator, whatever 3^6 is

OpenStudy (anonymous):

okay haha thanks !complicated but thank you

OpenStudy (anonymous):

You're Welcome